A project currently generates sales of $6 million, variable costs equal 40% of sales, and fixed costs are   $1.2 million. The firm’s tax rate is 30%. Assume all sales and expenses are cash item.

  1. What are the effects on cash flow, if sales increase from $6 million to $6.6 million?
    Cash flow_____by______?
  2. What are the effects of cash flow, if variable cost increase to 55% of sales?
    Cash flow_____by______?
Answer #1

a.

cash flow at sales of 6 million

=(6-(6*40%)-1.2)*(1-30%)

=1.68 million

cash flow at sales of 6.6 million

=(6.6-(6.6*40%)-1.2)*(1-30%)

=1.932 million

if sales increase from $6 million to $6.6 million, Cash flow increase by=1.932-1.68=0.252 million or 0.25 million

b.

cash flow at variable costs equal 40% of sales

=(6-(6*40%)-1.2)*(1-30%)

=1.68 million

cash flow at variable costs equal 55% of sales

=(6-(6*55%)-1.2)*(1-30%)

=1.05 million

if sales increase from $6 million to $6.6 million, Cash flow decrease by=1.68-1.05=0.63 million

the above is answer..